Find two consecutive integers such that the sum of five times the smaller and three times the largers is 67
Answer by mananth(16949)

Find two consecutive integers such that the sum of five times the smaller and three times the largers is 67
...
let x, x+1 be the integers
5 times smaller = 5x
3 times bigger = 3(x+1)
..
5x+3(x+1)=67
5x+3x+3=67
8x+3=67
8x=67-3
8x=64
/8
x= 8
8,9 are the integers
...
m.ananth@hotmail.ca